Two satellites orbit the Earth at the same altitude in circular orbits. One satellite has a mass of 250 kg, and the other has a mass of 150 kg. The orbital speed of the larger satellite is  


1
Expert's answer
2021-04-28T09:54:19-0400

Considering a circular orbit with the center-of-mass distance of 8000 km, the orbital speed of the larger satellite is  


"v=\sqrt{\frac{G(M+m)}{r}}=7074\text{ m/s}."